OpenStudy (anonymous):

2x+3y=3 5x+7y=0 Find the slope of the first line and the second line.

OpenStudy (anonymous):

y=(3-2x)/3=1-2/3*x -> slope is 2/3 (angle is tan(2/3)) y=-5/7*x -> slope is -5/7 (angle is tan(-5/7))